OpenStudy (anonymous):

Simplify the complex rational expression by the method of your choice. 1/x-2/1-1/x-2

OpenStudy (anonymous):



OpenStudy (anonymous):

I got 1/x-3

OpenStudy (campbell_st):

well start with the denominator \[1 - \frac{1}{x -2}\] you need a common denominator... \[\frac{x -2}{x -2} - \frac{1}{x -2} \] can you simplify this...?

OpenStudy (anonymous):

the answer is what i typed

hartnn (hartnn):

and your answer 1/(x-3) is correct. if you need to, you can metion the restriction that x \(\ne 2\), because you cancelled x-2 from numerator and denominator.